# Maintaining Plasticity in Deep Continual Learning

### v1

@@ Abstract
Modern deep-learning systems are specialized to problem settings in which training occurs once and then never again, as opposed to continual-learning settings in which training occurs continually. If deep-learning systems are applied in a continual learning setting, then it is well known that they may fail catastrophically to remember earlier examples. More fundamental, but less well known, is that they may also lose their ability to adapt to new data, a phenomenon called loss of plasticity. We show loss of plasticity using the MNIST and ImageNet datasets repurposed for continual learning as sequences of tasks. In ImageNet, binary classification performance dropped from 89% correct on an early task down to 77%, or to about the level of a linear network, on the 2000th task. Such loss of plasticity occurred with a wide range of deep network architectures, optimizers, and activation functions, and was not eased by batch normalization or dropout. In our experiments, loss of plasticity was correlated with the proliferation of dead units, with very large weights, and more generally with a loss of unit diversity. Loss of plasticity was substantially eased by $L^{2}$-regularization, particularly when combined with weight perturbation (Shrink and Perturb). We show that plasticity can be fully maintained by a new algorithm—called continual backpropagation—which is just like conventional backpropagation except that a small fraction of less-used units are reinitialized after each example. This continual injection of diversity appears to maintain plasticity indefinitely.  [^1] Code is available at https://github.com/shibhansh/loss-of-plasticity
@@ 1 Loss of Plasticity
Modern deep-learning systems have become specialized to problem settings in which training occurs once on a large data set and then never again. Such train-once settings were used in all the early successes of deep learning in speech recognition and image classification. Later, when deep learning was adapted to reinforcement learning (e.g., Mnih et al., 2015), techniques such as replay buffers and batching were introduced to make it very nearly a train-once setting. Recent applications of deep learning such as GPT-3 (Brown et al., 2020), and DallE (Ramesh et al., 2021) were also trained on data as a single large batch. Of course, in many applications the data distribution changes over time and training must continue in some form; the most common strategy in these cases has been to continually collect data and then, on occasion, train a new network from scratch, again in a train-once setting. The train-once setting has been integral to the design and practice of modern deep-learning methods.
@@
Deep learning systems, when exposed to new data, tend to forget most of what they have previously learned, a phenomenon known as “catastrophic forgetting.” In other words, deep learning methods do not maintain stability in continual learning problems. This phenomenon was first shown in early neural networks in the late 1900s (McCloskey & Cohen, 1989; French, 1999). Recently, with the advent of deep learning, catastrophic forgetting has received renewed attention as many papers have been dedicated to maintaining stability in deep continual learning (Kirkpatrick et al., 2017; Yoon et al., 2018; Aljundi et al., 2019; Golkar et al., 2019; Rajasegaran et al., 2019; Javed & White, 2019).
@@
Do deep-learning systems lose plasticity in continual learning problems? Some evidence that they do comes from the psychology literature of the early 2000s. Ellis & Lambon Ralph (2000), Zevin & Seidenberg (2002), and Bonin et al. (2004) showed loss of plasticity in early neural networks in supervised regression problems. These papers used a setup where a set of examples was presented to the network for a certain number of epochs, and then the training set was expanded with additional examples and training proceeded for another number of epochs. They found that the error for the examples in the original training set was lower than for the later-added examples after controlling for the number of epochs. These papers provided evidence that deep learning and the backpropagation algorithm (Rumelhart et al., 1986) on which it is founded tend to produce loss of plasticity. A limitation of these early works is that the networks used were relatively shallow by today’s standards, and the algorithms used are not those that are most popular today. For example, they did not use the ReLU activation function or the Adam optimizer. The early psychology research on artificial neural networks provides tantalizing hints, but does not provide a clear answer to the question of whether or not modern deep learning networks exhibit loss of plasticity.
@@
Turning now to the machine learning literature, we can find several recent studies suggesting loss of plasticity in deep learning. Chaudhry et al. (2018) observed loss of plasticity in continual image classification problems. Their experiments are not fully satisfactory as a demonstration of loss of plasticity because they introduced other confounding variables. In their setup, when a new task was presented, new outputs, called heads, were added to the network; the number of outputs grew as additional tasks were encountered. Loss of plasticity was thereby confounded with the effects of interference from old heads. Chaudhry et al. found that when old heads were removed at the start of a new task, the loss of plasticity was minimal, which suggests that the loss of plasticity they observed was mainly due to interference from old heads. A second limitation of Chaudhry et al.’s study is that it used only ten tasks and thus did not assess loss of plasticity when deep learning methods face a long sequence of tasks.
@@
Ash and Adams (2020) showed that initializing a neural network by first training it on a fraction of the dataset can lead to much lower final performance compared to the case in which the network is trained once on the entire dataset. The worse performance after pretraining is an important example of loss of plasticity in deep learning and hints at the possibility of a major weakness in full continual learning with many task changes or non-stationarity. Berariu et al. (2021) extended Ash and Adams’ work to show that the more pretraining stages there were, the worse the final performance. However, Berariu et al.’s experiments were performed in a stationary problem setting, and the loss of plasticity they observed was small. Although the results in these papers hint at a fundamental loss of plasticity in deep learning systems, none directly demonstrated loss of plasticity in continual learning.
@@
More evidence for loss of plasticity in modern deep learning has been obtained in the reinforcement learning community, where recent papers have shown substantial loss of plasticity. Nishikin et al. (2022) showed that early learning in reinforcement learning problems could harm later learning, an effect they termed “primacy bias.” This result could be due to the loss of plasticity of deep learning networks in continual learning settings, as reinforcement learning is inherently continual due to changes in the policy. Lyle et al. (2022) also showed that some deep reinforcement learning agents could lose the ability to learn new functions over time. These are important data points, but the conclusions that can be drawn from them are limited by the inherent complexity of modern deep reinforcement learning. All of these papers, both from the psychology literature at the turn of the century and recent papers in machine learning and reinforcement learning, are evidence that deep-learning systems lose plasticity, but fall short of a fully satisfactory demonstration of the phenomena.
@@
After showing the severity and generality of loss of plasticity in deep learning, we seek a deeper under-standing of its causes. This leads us to explore existing methods that might ameliorate loss of plasticity. We find that some methods, like Adam (Kingma & Ba, 2015) and dropout (Hinton et al., 2012), make the loss of plasticity significantly worse, while other methods, like $L^{2}$ regularization (Goodfellow et al., 2016, pp. 227-230) and Shrink and Perturb (Ash & Adams, 2020), substantially ease the loss of plasticity in many cases. Finally, we propose a new algorithm—continual backpropagation—that robustly solves the problem of loss of plasticity in both our systematic experiments with supervised learning problems and in a preliminary experiment with a reinforcement learning problem. The usual backpropagation algorithm has two crucial components: 1) initialization of the network’s connection weights with small random numbers and 2) stochastic gradient descent on every presentation of a training example (Rumelhart, Hinton & Williams, 1986). Continual backpropagation extends the initialization to all time steps, by selectively reinitializing a small fraction of low-utility hidden units at every presentation of a training example. Continual backpropagation performs both gradient descent and selective reinitialization continually.
@@ 2 Loss of Plasticity in ImageNet
Imagenet is a large database of images and their labels that has been influential throughout machine learning (Deng et al., 2009; Image-net.org) and that played a key role in the rise of deep learning (Krizhevsky et al., 2012). ImageNet allowed researchers to conclusively show that deep learning can solve a problem like object recognition, a hallmark of intelligence. Classical machine learning methods had a classification error rate of 25% (Russakovsky et al., 2015) when deep-learning methods reduced it to below 5% (Hu et al., 2018), the human error rate on the dataset.
The ImageNet database comprises millions of images labelled by nouns (classes) such as animals and everyday objects. In the standard train-once setting, the dataset is partitioned into a training set and a test set. The learning system is first trained on the training set, and then the learner is frozen and its performance is measured on the test set.
We sought to adapt Imagenet from the train-once setting to the continual-learning setting while mini-mizing all other changes. The Imagenet database we used consisted of 1000 classes, each of 700 images. We constructed from these a near-endless sequence of binary classification tasks. For example, the first task might be to distinguish Class1 from Class2, and the second task might be to distinguish Class3 from Class4. Each binary classification task was handled in conventional way. The 700 images for each class were divided into 600 images for a training set and 100 images for a test set. On each task, the deep learning network was first trained on the training set of 1200 images, and then its classification accuracy was assessed on the test set of 200 images. That training consisted of multiple passes through the training set, called epochs. After training and test on one task, the next task was begun based on a different pair of classes. By selecting the pairs of tasks randomly without replacement, we generated a sequence of 2000 different tasks. All tasks used the downsampled 32x32 version of the ImageNet dataset, as is often done to save computation (Chrabaszcz et al., 2017). We call the resulting continual-learning problem Continual ImageNet.
@@
We applied a wide variety of standard deep-learning networks to Continual ImageNet. We describe here results with a specific, representative network, using convolutional layers and the ReLU activation function. The network had three convolutional-plus-max-pooling layers followed by three fully connected layers, as detailed in Table 1. This network is narrower than those commonly used on the ImageNet dataset because here we are trying to discriminate only two classes at a time rather than all 1000. The final layer consisted of just two units, the heads, corresponding to the two classes. At task changes, the input weights of the heads were reset to zero. Resetting the heads in this way can be viewed as introducing new heads for the new classes. This resetting of the output weights is not ideal for studying plasticity, as the learning system gets access to privileged information on the timing of task changes (and we do not use it in the experiments reported later in this paper). We use it here because it is closest to the standard practice in deep continual learning, which is to introduce new heads for new classes.
We tested many variations of the network architecture and hyper-parameters to obtain good performance on the first binary classification task. The networks were trained using stochastic gradient descent (SGD) with momentum on the cross-entropy loss and initialized once prior to the first task. For each task, all learning algorithms performed 250 passes through the training set using mini-batches of size 100. The momentum hyperparameter was 0.9. We tested various step-size parameters but are only presenting the performance for step sizes 0.01, 0.001, and 0.0001 for clarity of the Figure. We performed 30 runs for each hyperparameter value, varying the sequence of tasks and other randomness. Across different hyperparameters, the same sequences of pairs of classes were used.
@@
The largest step size, 0.01, performed the best on the first two tasks, but performance then fell on subsequent tasks, eventually reaching a level below the linear baseline. For even larger step sizes like 0.1, the network diverged after the first few tasks. At the smaller step sizes, performance rose initially and then fell and was only slightly better than the linear baseline after 2000 tasks. The performance for intermediate step-sizes like 0.003 followed the general trend of the ones presented in Figure 1. We have found this to be a common pattern in our experiments: for a well-tuned network, performance first improves, then falls substantially, ending near or below the linear baseline. We have observed this pattern for many network architectures, hyper-parameters, and optimizers. The specific choices of network architecture, hyper-parameters, and optimizers affect when the performance starts to drop, but we observed a severe performance drop for a wide range of choices.
This experiment was designed to closely follow standard deep-learning practice in every way possible, except in requiring the network to keep learning. Failure to learn better than a linear network in later tasks is substantial evidence that the standard practice of deep learning simply doesn’t work in continual problems. However, there are of course many other variations of deep learning that could be tried. To reach the next level of confidence and understanding we switch in the next section to a less computationally complex problem where we can study the phenomena more thoroughly.
@@ 3 Robust Loss of Plasticity in Permuted MNIST
MNIST (LeCun & Cortes, 1998) is one of the most common supervised-learning datasets used in deep learning. It consists of 60,000 28x28 grayscale images of the handwitten digits from 0 to 9 together with their correct labels. For example, the left image in Figure 2 shows an image that is labelled by the digit 7. The smaller number of classes and the simpler images enables much smaller deep learning networks to perform well on this dataset than are needed on ImageNet. The smaller networks in turn mean much less computation is needed to perform the experiments and thus many more and more varied experiments can be done, enabling us to perform deeper and more extensive studies of plasticity. As with ImageNet, MNIST is typically used in a train-once setting.
Figure 2: Left: An MNIST image with the label ‘7’; Right: A corresponding permuted image.
We created a continual supervised learning problem using permuted MNIST datasets (Goodfellow et al., 2014; Zenke et al., 2017). An individual permuted MNIST dataset is created by permuting the pixels in the original MNIST dataset. For example, the pixel that originally appeared in the upper-left corner might be moved to the lower-right corner, or to any other particular place in the image, and so on for all the other pixels. The right image in Figure 2 is an example of such a permuted image for an image labelled with the digit 7. Given a way of permuting, all 60,000 images are permuted in the same way to produce the new permuted MNIST dataset.
@@
We applied feed-forward neural networks with three hidden layers to Online Permuted MNIST. We did not use convolutional layers, as they could not be helpful on the permuted problem (in MNIST, convolutional layers are often not used even on the standard, non-permuted problem). For each example, the network estimated the probabilities of each of the 10 classes (0–9), compared them to the correct label, and performed SGD on the cross-entropy loss. As a measure of performance, we recorded the percentage of times the network’s highest probability label was the correct over a tasks 60,000 images. We plot this per-task performance measure versus task number in the figures below. The weights were initialized according to a Kaiming distribution.
@@
The first panel of Figure 3 shows the progression of performance across tasks for a network of with 2000 units per layer, and various values of the step-size parameter. Note that performance first increased across tasks, then began falling steadily across all subsequent tasks. These results are consistent with the loss of plasticity that we observed in ImageNet.
Figure 3: Loss of plasticity in Online Permuted MNIST is robust over step sizes, network sizes, and rates of change.
@@
Next we varied the network size. Instead of 2000 units per layer, we tried 100, 1000, and 10,000 units per layer. In this experiment we ran for only 150 tasks, primarily because the largest network took much longer to run. The time courses of performance at good step sizes for each network size are shown in the middle panel of Figure 3. The loss of plasticity with continued training is most pronounced at the smaller network sizes, but even the largest networks showed some loss of plasticity.

### v3

@@ Abstract
Modern deep-learning systems are specialized to problem settings in which training occurs once and then never again, as opposed to continual-learning settings in which training occurs continually. If deep-learning systems are applied in a continual learning setting, then it is well known that they may fail to remember earlier examples. More fundamental, but less well known, is that they may also lose their ability to learn on new examples, a phenomenon called loss of plasticity. We provide direct demonstrations of loss of plasticity using the MNIST and ImageNet datasets repurposed for continual learning as sequences of tasks. In ImageNet, binary classification performance dropped from 89% accuracy on an early task down to 77%, about the level of a linear network, on the 2000th task. Loss of plasticity occurred with a wide range of deep network architectures, optimizers, activation functions, batch normalization, dropout, but was substantially eased by $L^{2}$-regularization, particularly when combined with weight perturbation. Further, we introduce a new algorithm—continual backpropagation—which slightly modifies conventional backpropagation to reinitialize a small fraction of less-used units after each example and appears to maintain plasticity indefinitely.  [^1] Code is available at https://github.com/shibhansh/loss-of-plasticity
@@ 1 Loss of Plasticity
Modern deep-learning systems have become specialized to problem settings in which training occurs once on a large data set and then never again. Such train-once settings were used in all the early successes of deep learning in speech recognition hinton2012speech and image classification krishevsky2012 . Later, when deep learning was adapted to reinforcement learning (e.g., Mnih et al. mnih2015human ), techniques such as replay buffers and batching were introduced to make it very nearly a train-once setting. Recent applications of deep learning such as GPT-3 brown2020 and DallE dallE were also trained on data as a single large batch. Of course, in many applications the data distribution changes over time and training must continue in some form; the most common strategy in these cases has been to continually collect data and then, on occasion, train a new network from scratch, again in a train-once setting. The train-once setting has been integral to the design and practice of modern deep-learning methods.
@@
Deep learning systems, when exposed to new data, tend to forget most of what they have previously learned, a phenomenon known as “catastrophic forgetting.” In other words, deep learning methods do not maintain stability in continual learning problems. This phenomenon was first shown in early neural networks in the late 1900s mccloskey1989 ; french1999 . Recently, with the advent of deep learning, catastrophic forgetting has received renewed attention as many papers have been dedicated to maintaining stability in deep continual learning kirkpatrick2017 ; yoon2018 ; aljundi2019 ; golkar2019 ; riemer2019 ; rajasegaran2019 ; javed2019 ; aspects-of-cl .
@@
Do deep-learning systems lose plasticity in continual learning problems? Some evidence that they do comes from the psychology literature of the early 2000s. Ellis and Lambon Ralph ellis2000age , Zevin and Seidenberg, zevin2002 , and Bonin et al. bonin2004 showed loss of plasticity in early neural networks in supervised regression problems. These papers used a setup where a set of examples was presented to the network for a certain number of epochs, and then the training set was expanded with additional examples and training proceeded for another number of epochs. They found that the error for the examples in the original training set was lower than for the later-added examples after controlling for the number of epochs. These papers provided evidence that deep learning and the backpropagation algorithm rumelhart1986learning on which it is founded tend to produce loss of plasticity. A limitation of these early works is that the networks used were relatively shallow by today’s standards, and the algorithms used are not those that are most popular today. For example, they did not use the ReLU activation function or the Adam optimizer. The early psychology research on artificial neural networks provides tantalizing hints, but does not provide a clear answer to the question of whether or not modern deep learning networks exhibit loss of plasticity.
@@
Turning now to the machine learning literature, we can find several recent studies suggesting loss of plasticity in deep learning. Chaudhry et al. intransigence2018 observed loss of plasticity in continual image classification problems. Their experiments are not fully satisfactory as a demonstration of loss of plasticity because they introduced other confounding variables. In their setup, when a new task was presented, new outputs, called heads, were added to the network; the number of outputs grew as additional tasks were encountered. Loss of plasticity was thereby confounded with the effects of interference from old heads. Chaudhry et al. found that when old heads were removed at the start of a new task, the loss of plasticity was minimal, which suggests that the loss of plasticity they observed was mainly due to interference from old heads. A second limitation of Chaudhry et al.’s study is that it used only ten tasks and thus did not assess loss of plasticity when deep learning methods face a long sequence of tasks.
@@
Ash and Adams WarmStart2020 showed that initializing a neural network by first training it on a fraction of the dataset can lead to much lower final performance compared to the case in which the network is trained once on the entire dataset. The worse performance after pretraining is an important example of loss of plasticity in deep learning and hints at the possibility of a major weakness in full continual learning with many task changes or non-stationarity. Berariu et al. plasticitystudy2021 extended Ash and Adams’ work to show that the more pretraining stages there were, the worse the final performance. However, Berariu et al.’s experiments were performed in a stationary problem setting, and the loss of plasticity they observed was small. Although the results in these papers hint at a fundamental loss of plasticity in deep learning systems, none directly demonstrated loss of plasticity in continual learning.
@@
More evidence for loss of plasticity in modern deep learning has been obtained in the reinforcement learning community, where recent papers have shown substantial loss of plasticity. Nishikin et al. primacy2022 showed that early learning in reinforcement learning problems could harm later learning, an effect they termed “primacy bias.” This result could be due to the loss of plasticity of deep learning networks in continual learning settings, as reinforcement learning is inherently continual due to changes in the policy. Lyle et al. capacityloss2022 also showed that some deep reinforcement learning agents could lose the ability to learn new functions over time. These are important data points, but the conclusions that can be drawn from them are limited by the inherent complexity of modern deep reinforcement learning. All of these papers, both from the psychology literature at the turn of the century and recent papers in machine learning and reinforcement learning, are evidence that deep-learning systems lose plasticity, but fall short of a fully satisfactory demonstration of the phenomena.
@@
After showing the severity and generality of loss of plasticity in deep learning, we seek a deeper under-standing of its causes. This leads us to explore existing methods that might ameliorate loss of plasticity. We find that some methods, like Adam kingma2015 and dropout Hinton2012Dropout1 , significantly exacerbate loss of plasticity, while other methods, like $L^{2}$ regularization (Goodfellow-et-al-2016 , pp. 227–230) and Shrink and Perturb WarmStart2020 , substantially ease the loss of plasticity in many cases. Finally, we propose a new algorithm—continual backpropagation—that robustly solves the problem of loss of plasticity in both our systematic experiments with supervised learning problems and in a preliminary experiment with a reinforcement learning problem. The usual backpropagation algorithm has two crucial components: 1) initialization of the network’s connection weights with small random numbers and 2) stochastic gradient descent on every presentation of a training example rumelhart1986learning . Continual backpropagation extends the initialization to all time steps, by selectively reinitializing a small fraction of low-utility hidden units at every presentation of a training example. Continual back-propagation performs both gradient descent and selective reinitialization continually.
@@ 2 Loss of Plasticity in ImageNet
The main results of this paper are contained in this section. Recall that the primary point of this paper is to establish in a more definitive way that deep learning methods lose the ability to learn in continual learning. To establish this, we use the classical test problem known as ImageNet and adapt it to continual learning. ImageNet has been historically important for deep learning and is still a widely used test bed. A classical deep learning problem like ImageNet is the best place to show the issue of loss of plasticity. In this section, we show that deep learning suffers from loss of plasticity in a continual learning problem based on ImageNet.
Imagenet is a large database of images and their labels that has been influential throughout machine learning (deng2009 , Image-net.org) and that played a pivotal role in the rise of deep learning krishevsky2012 . ImageNet allowed researchers to conclusively show that deep learning can solve a problem like object recognition, a hallmark of intelligence. Classical machine learning methods had a classification error rate of 25% russakovsky2015imagenet when deep-learning methods reduced it to below 5% hu2018squeeze , the human error rate on the dataset.
The ImageNet database comprises millions of images labelled by nouns (classes) such as types of animals and everyday objects. In the standard train-once setting, the dataset is partitioned into a training set and a test set. The learning system is first trained on the training set, and then the learner is frozen, and its performance is measured on the test set.
@@
We sought to adapt Imagenet from the train-once setting to the continual-learning setting while mini-mizing all other changes. The Imagenet database we used consists of 1000 classes, each of 700 images. We constructed from these a near-endless sequence of binary classification tasks. For example, the first task might be to distinguish Class1 from Class2, and the second task might be to distinguish Class3 from Class4. Each binary classification task was handled in conventional way. The 700 images for each class were divided into 600 images for a training set and 100 images for a test set. On each task, the deep learning network was first trained on the training set of 1200 images, and then its classification accuracy was assessed on the test set of 200 images. That training consisted of multiple passes through the training set, called epochs. After training and testing on one task, the next task was begun based on a different pair of classes. By selecting the pairs of tasks randomly without replacement, we generated a sequence of 2000 different tasks. All tasks used the downsampled 32x32 version of the ImageNet dataset, as is often done to save computation ImagenetDownsampled . We call the resulting continual-learning problem Continual ImageNet. In this section, we use Continual ImageNet to establish loss of plasticity.
To show that deep learning loses the ability to learn, we applied a wide variety of standard deep learning networks to Continual ImageNet. We describe here a specific, representative convolutional network with ReLU activation function. The network had three convolutional-plus-max-pooling layers followed by three fully connected layers, as detailed in Table 1. This network is narrower than those commonly used on the ImageNet dataset because here we are trying to discriminate only two classes at a time rather than all 1000. The final layer consisted of just two units, the heads, corresponding to the two classes. At task changes, the input weights of the heads were reset to zero. Resetting the heads in this way can be viewed as introducing new heads for the new classes. This resetting of the output weights is not ideal for studying plasticity, as the learning system gets access to privileged information on the timing of task changes (and we do not use it in the experiments reported later in this paper). We use it here because it is closest to the standard practice in deep continual learning, which is to introduce new heads for new classes kirkpatrick2017 ; yoon2018 ; aljundi2019 ; golkar2019 .
@@
We tested many variations of the network architectures, hyperparameters, and optimizers to obtain good performance on the first binary classification task. We now describe the details of one such training procedure applied to the network described in Table 1. The network was trained using stochastic gradient descent (SGD) with momentum on the cross-entropy loss and initialized once prior to the first task. For each task, the learning algorithm performed 250 passes through the training set using mini-batches of size 100. The momentum hyperparameter was 0.9. We tested various step-size parameters but are only presenting the performance for step sizes 0.01, 0.001, and 0.0001 for clarity of the figure. We performed 30 runs for each hyperparameter value, varying the sequence of tasks and other randomness. Across different hyperparameters, the same sequences of pairs of classes were used.
@@
The network had poor performance on the 2000th task for all hyperparameter values. The largest step size, 0.01, performed the best on the first two tasks, but performance then fell on subsequent tasks, eventually reaching a level below the linear baseline. For even larger step sizes like 0.1, the network diverged after the first few tasks. At the smaller step sizes, performance rose initially and then fell and was only slightly better than the linear baseline after 2000 tasks. The performance for intermediate step sizes like 0.003 followed the general trend of the step sizes presented in Figure 1. We have found this to be a common pattern in our experiments: for a well-tuned network, performance first improves, then falls substantially, ending near or below the linear baseline. We have observed this pattern for many network architectures, hyperparameters, and optimizers. The specific choices of network architecture, hyperparameters, and optimizers affect when the performance starts to drop, but we observed a severe performance drop for a wide range of choices. To increase confidence in the results and to ensure that the performance drop is not due to some bug in the implementation, one of the authors independently reproduced the results of this experiment.
This experiment was designed to closely follow standard deep-learning practice in every way possible, except in requiring the network to keep learning. Failure to learn better than a linear network in later tasks is substantial evidence that the standard practice of deep learning simply does not work in continual problems. We showed systematically through a direct experiment that deep learning consistently loses the ability to learn new things, i.e. deep learning loses plasticity, in continual learning problem. However, there are of course many other variations of deep learning that could be tried. To reach the next level of confidence and understanding we switch in the next section to a less computationally complex problem where we can study the nuances of the phenomena more thoroughly.
@@ 3 Robust Loss of Plasticity in Permuted MNIST
We now use a computationally cheap problem based on the MNIST dataset lecun-mnist-1998 to test the generality of loss of plasticity. MNIST is one of the most common supervised-learning datasets used in deep learning. It consists of 60,000 28x28 grayscale images of the handwitten digits from 0 to 9 together with their correct labels. For example, the left image in Figure 2a shows an image that is labelled by the digit 7. The smaller number of classes and the simpler images enables much smaller deep learning networks to perform well on this dataset than are needed on ImageNet. The smaller networks in turn mean much less computation is needed to perform the experiments and thus experiments can be performed in greater quantities and under a variety of different conditions, enabling us to perform deeper and more extensive studies of plasticity. As with ImageNet, MNIST is typically used in a train-once setting.
Figure 2: a: Left: An MNIST image with the label ‘7’; Right: A corresponding permuted image. b: Loss of plasticity in Online Permuted MNIST is robust over step sizes, network sizes, and rates of change.
We created a continual supervised learning problem using permuted MNIST datasets goodfellow2014 ; zenke2017 . An individual permuted MNIST dataset is created by permuting the pixels in the original MNIST dataset. For example, the pixel that originally appeared in the upper-left corner might be moved to the lower-right corner, or to any other particular place in the image, and so on for all the other pixels. The right image in Figure 2a is an example of such a permuted image for an image labelled with the digit 7. Given a way of permuting, all 60,000 images are permuted in the same way to produce the new permuted MNIST dataset.
@@
We applied feed-forward neural networks with three hidden layers to Online Permuted MNIST. We did not use convolutional layers, as they could not be helpful on the permuted problem because the spatial information is lost (in MNIST, convolutional layers are often not used even on the standard, non-permuted problem). For each example, the network estimated the probabilities of each of the 10 classes (0–9), compared them to the correct label, and performed SGD on the cross-entropy loss. As a measure of performance, we recorded the percentage of times the network’s highest probability label was the correct one over a task’s 60,000 images. We plot this per-task performance measure versus task number in Figure 2b. The weights were initialized according to a Kaiming distribution.
@@
The first panel of Figure 2b shows the progression of performance across tasks for a network of with 2000 units per layer, and various values of the step-size parameter. Note that performance first increased across tasks, then began falling steadily across all subsequent tasks. This drop in performance means that the network is slowly losing the ability to learn on new tasks. This loss of plasticiy is consistent with the loss of plasticity that we observed in ImageNet.
@@
Next we varied the network size. Instead of 2000 units per layer, we tried 100, 1000, and 10,000 units per layer. In this experiment we ran for only 150 tasks, primarily because the largest network took much longer to run. The time courses of performance at good step sizes for each network size are shown in the middle panel of Figure 2b. The loss of plasticity with continued training is most pronounced at the smaller network sizes, but even the largest networks show some loss of plasticity.